A battery powered toy car accelerates from 0.0 m/s to 8.5 m/s. It’s mass is 0.55kg. The chemical store in its battery is decreased by 30J. Calculate the efficiency of this process


1
Expert's answer
2021-01-12T12:12:12-0500
"\eta=\frac{K}{E}\\\eta=\frac{mv^2}{2E}\\\eta=\frac{0.55\cdot8.5^2}{2\cdot30}=0.66"